OpenStudy (anonymous):

find the derivative at 12 of the inverse of the function f[x] = .25(x^5 + 2x^3)

OpenStudy (anonymous):

I know so far that i use 1/f'[f^-1[x]]. However, I do not know how to apply this

OpenStudy (anonymous):

Did you say something tkhunny?

OpenStudy (tkhunny):

You have f(x). Find f'(x).

OpenStudy (anonymous):

ok f'[x] = .25(6x^2 + 5x^4) Now what?

OpenStudy (tkhunny):

Well, there is little hope if finding \(f^{-1}(x)\) explicitly, so this must be a numerical problem. This is fine because we have x = 12 as the place where we are interested.

OpenStudy (anonymous):

I tried using the definition of an inverse (domain and range flipped), but there is NO way to solve that equation. I am desperate

OpenStudy (tkhunny):

You do not need to solve for the inverse. You need to evaluate the inverse at x = 12. f(2) = 12 -- Serendipity provides that this is the ONLY Real Solution, so there is no ambiguity. \(f^{-1}(12) = 2\)

OpenStudy (tkhunny):

So, \(f'(f^{-1}(12)) = f'(2)\). Can you take it from there?

OpenStudy (anonymous):

Yes I can easily take it from there. However, can you elaborate on how you got 2? I don't see any hint in the problem that would let us find 2 in there

OpenStudy (tkhunny):

Well, somehow, you just have to solve for it. \(\dfrac{1}{4}\left(x^{5}+2x^{3}\right) = 12\) \(\left(x^{5}+2x^{3}\right) = 48\) \(x^{5}+2x^{3} - 48 = 0\) The HAS a positive solution in the Real Numbers (do you know why?). If it's Rational, which I hope it is, it will be a factor of 48 (do you know why?). This narrows the choices considerably.

OpenStudy (anonymous):

i'm not sure why we know why it has a real positive solution, but I do know why it has a factor of 48. Can you elaborate on that?

OpenStudy (tkhunny):

Perhaps you have heard of "Descartes' Rule of Signs"?